# Approaching the Quantum Speed Limit in Quantum Gates with Geometric Control

### v1

@@ Abstract
We present a geometric optimization method for implementing quantum gates by optimally controlling the Hamiltonian parameters, aiming to approach the Mandelstam-Tamm Quantum Speed Limit (MT-QSL). Achieving this bound requires satisfying precise geometric conditions governing the evolution of quantum states. We extend this geometric framework to quantum unitary operators in arbitrary dimensions and analyze the conditions necessary for saturation of the bound. Additionally, we demonstrate that the quantum brachistochrone, when generalized to operators, does not generally reach the MT-QSL bound. Finally, we propose a systematic optimal control strategy based on geometric principles to approach the quantum speed limit for unitary driving.
@@
Approaching optimality in quantum driving is both a challenging and crucial task for quantum information systems. Reducing evolution time while maintaining constant resources enhances efficiency and can potentially increase the resilience of quantum protocols against external perturbations. The concept of the Quantum Speed Limit (QSL), which defines the minimal evolution time required for an initial state to reach an orthogonal subspace, has been a subject of growing interest and extensive investigation [1, 2, 3, 4, 5, 6, 7, 8], with the recent generalization to open  [9, 10, 11, 12], non-Markovian [13] and non-linear [14, 15] quantum systems. Among recent developments, we highlight the search for unified geometric definitions of the QSL [16], the investigation of the relation of the QSL with the system quantumness [17, 18], with interference effects [19, 20], or with statistical properties [21, 22] of the system, and finally an effort to approach the QSL in concrete experimental setups [23]. The QSL is deeply linked to the geometry of quantum state space [24] and involves identifying quantum trajectories that correspond to geodesics in the Fubini-Study metric. When formulated as a quantum brachistochrone problem, it allows for analytical [25, 26] or approximate [27] solutions for specific classes of Hamiltonians. However, optimal solutions to the quantum brachistochrone problem do not always saturate the QSL bounds associated with the evolution of a single quantum state [28]. While a significant body of literature has focused on the QSL for the evolution of individual states, the exploration of QSLs governing the evolution of sets of quantum states remains in its early stages [29, 30, 31, 32].
@@
The goal of this paper is to present an alternative derivation of such bounds, inspired by the approach of L. Vaidman [3], which also introduces a geometric optimization method for implementing time-optimal unitary quantum gates. More specifically, we identify two necessary criteria for optimal driving: one involves a phase condition, while the other has a geometric interpretation. Indeed, the deviation from the optimal Mandelstam-Tamm QSL bound increases as the time evolution of the operator deviates from a specific time-dependent plane in the Hilbert space of quantum operators.
@@
In Sec. I, we generalize the Mandelstam-Tamm bound on the quantum speed limit using an approach similar to that developed by L. Vaidman [3]. We then discuss the generalization of the Margolus-Levitin bound for time evolution operators. In Sec. II, we explore the conditions under which the bound is saturated, illustrating these with various examples, and analyze the quantum brachistochrone from the QSL perspective. Finally, in Sec. III, we use the geometric framework to identify the additional constraints that must be incorporated into an optimal control approach to achieve the maximal Quantum Speed Limit (QSL) for a given unitary operator. As intuitively expected, the efficiency of optimized quantum driving relative to the QSL is shown to depend on the degree of control over the Hamiltonian parameters. To illustrate this, we consider two scenarios using a concrete example of unitary gates in a four-dimensional space: one with full control over the Hamiltonian parameters, and another with more restricted control.
@@ I Quantum Speed Limits for unitary operators
In this paragraph, we derive the two standard bounds for the quantum speed limit in the context of quantum control of time evolution operators. We consider a $d$-dimensional Hilbert space $\mathcal{H}$ and the associated space of linear operators $\mathcal{L}(\mathcal{H})$, on which we define the scalar product $\langle U_{1},U_{2}\rangle=\frac{1}{d}{\rm Tr}\left[U_{1}^{\dagger}U_{2}\right]$. Unitary operators $U$ have unit norm $\langle U,U\rangle=1$.
@@ I.1 Mandelstamm and Tamm bound for quantum operators
Using geometric arguments, L. Vaidman derived a minimal time required to drive a given initial quantum state to any orthogonal quantum state  [3]. Following a similar approach, we derive a QSL for the full quantum evolution operator and establish a lower bound on the evolution time required to reach an orthogonal quantum operator.
@@
where $U_{\perp}(t)$ is a linear operator such that $\langle U(t),U_{\perp}(t)\rangle=0$ and $\langle U_{\perp}(t),U_{\perp}(t)\rangle=1$ (note that $U_{\perp}(t)$ is in general non-unitary). By projecting both sides of the equation onto the operator $U(t)$, we obtain $\lambda_{/\!/}(t)=\frac{1}{d}{\rm Tr}[U^{\dagger}(t)H(t)U(t)]=\frac{1}{d}{\rm Tr}[H(t)]=\langle H(t)\rangle$ where we define $\langle O\rangle=\frac{1}{d}{\rm Tr}[O]$ for any operator $O$. Next, we take the scalar product with the Hermitian conjugate $H(t)U(t)$ on both sides of Eq. (5), and readily obtain $\frac{1}{d}{\rm Tr}[H(t)^{2}]=|\lambda_{/\!/}(t)|^{2}+|\lambda_{\perp}(t)|^{2}$ where we have used $H(t)^{\dagger}=H(t)$, $\langle U(t),U(t)\rangle=\langle U_{\perp}(t),U_{\perp}(t)\rangle=1$ and $\langle U_{\perp}(t),U(t)\rangle=\langle U(t),U_{\perp}(t)\rangle=0.$ As a result, we find $|\lambda_{\perp}(t)|^{2}=\langle H(t)^{2}\rangle-\langle H(t)\rangle^{2}$, i.e. $|\lambda_{\perp}(t)|=\Delta H(t).$ Consequently, the operator $U_{\perp}$ can be expressed as
$U_{\perp}(t)=\frac{1}{\Delta H(t)}\left(H(t).U(t)-\langle H(t)\rangle U(t)\right)$ (2)
@@
$\frac{d|\langle U_{0},U(t)\rangle|^{2}}{dt}=\frac{-2\Delta H(t)}{\hbar}{\rm Re}\left[(-i)\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle\right]$
We have omitted the term proportional to the average energy, which corresponds to a purely imaginary number. The initial operator $U_{0}$ can be decomposed as
@@
$|\langle U_{0},U_{\perp}(t)\rangle|$ $=$ $\sqrt{1-|\langle U_{0},U(t)\rangle|^{2}-||U_{\perp\perp}(t)||^{2}}$ (6)
@@
$\frac{d|\langle U_{0},U(t)\rangle|^{2}}{dt}\leq\frac{2\Delta H(t)}{\hbar}|\langle U_{0},U(t)\rangle|\sqrt{1-|\langle U_{0},U(t)\rangle|^{2}}$ (8)
By setting $\cos\phi=|\langle U_{0},U(t)\rangle|$, the inequality (8) can expressed as an upper bound on the time derivative of the angle $\phi$:
$|\dot{\phi}|\leq\frac{\Delta H}{\hbar}$ (9)
To saturate this QSL, two conditions must be simultaneously fulfilled
$|\langle U_{0},U(t)\rangle|^{2}+|\langle U_{0},U_{\perp}(t)\rangle|^{2}$ $=$ 11 11
${\rm Re}\left[\langle U(t),U_{0}\rangle\langle U_{0}(t),U_{\perp}(t)\rangle\right]$ $=$ 00 0 (10)
The first condition states that $U_{\perp\perp}(t)=0,$ meaning that the time derivative $\dot{U}(t)$ must always lie in the plane spanned by $\{U_{0},U(t)\}$. Since the evolution operator $U(t)$ is also required to remain unitary at all times, this imposes a significant constraint on the possible quantum paths that saturate the QSL bound. In practice, it is not guaranteed that such a path exists in Hilbert spaces with dimensions larger than two. In such cases, one can quantify the deviation from the QSL when these conditions are not satisfied. Setting $\alpha(t)=||U_{\perp\perp}(t)||^{2}=\sin^{2}\phi(t)-|\langle U_{0},U_{\perp}(t)\rangle|^{2}$ and $\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle=|\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle|e^{i\beta(t)}$, one has
$\dot{\phi}=\frac{\Delta H(t)}{\hbar}\sqrt{1-\frac{|\alpha(t)|^{2}}{\sin\phi(t)^{2}}}\sin\beta(t).$ (11)
The ideal case of QSL saturation corresponds to $\alpha(t)=0$ (plane condition) and to $\beta(t)=\pm\frac{\pi}{2}$ (phase condition) at all times. An optimization method based on these two relations will be presented in Sec. III.
@@ I.2 Margolus-Levitin quantum speed limit
Interestingly, one can also retrieve the Margolus-Levitin bound [33] in this operatorial formalism. For this purpose, we consider a time-independent Hamiltonian $H_{0}=\sum_{n=1}^{d}E_{n}|n\rangle\langle n|$ with $E_{1}\leq E_{2}\leq...\leq E_{d}$ with eventually degenerate eigenvalues. The associated evolution operator reads $U(t)=\sum_{n=1}^{d}e^{-\frac{i}{\hbar}E_{n}t}|n\rangle\langle n|,$ and one finds
$D(t)=\frac{1}{d}{\rm Tr}\left[U_{0}^{\dagger}U(t)\right]=\frac{1}{d}\sum_{n=1}^{d}e^{-\frac{i}{\hbar}E_{n}t}$ (12)
where we have considered the initial evolution operator $U_{0}=\mathbb{I}$ (the same argument could be applied to a different gate $U_{f}$). One can proceed in the same way as in the proof for single states (with here $|c_{n}|^{2}\equiv 1/d$). Thus, we have that
${\rm Re}\left[D(t)\right]$ $=$ $\frac{1}{d}\sum_{n=1}^{d}\cos\left(\frac{E_{n}}{\hbar}t\right)$ (13)
$\geq$ $\frac{1}{d}\sum_{n}\left[1-\frac{2}{\pi}\left(\frac{E_{n}t}{\hbar}+\sin\left(\frac{E_{n}t}{\hbar}\right)\right)\right]$
$\geq$ $1-\frac{2}{\pi}\langle H_{0}\rangle\frac{t}{\hbar}+\frac{2}{\pi}{\rm Im}[D(t)]$
@@
valid for $x\geq 0$, which becomes an equality if and only if $x=0,\pi$. Assuming that $D(\tau)=0$, i.e., that the evolution leads to an orthogonal gate after time $\tau$, the previous inequality gives a lower bound on this duration
$\tau\geq\frac{\pi\hbar}{2\langle H_{0}\rangle}$ (15)
This QSL bound is valid for any constant and positive Hamiltonian $H_{0}$. The replacement $H_{0}^{\prime}=H_{0}-E^{\prime}$ merely adds a global phase factor and thus leaves invariant the time $\tau$. As long as $E^{\prime}\leq E_{1}$ to guarantee the positivity of the Hamiltonian, the argument above yields a similar bound with $\langle H_{0}^{\prime}\rangle=\langle H_{0}\rangle-E^{\prime}$ instead of $\langle H_{0}\rangle.$ The tighter bound is obtained with $E^{\prime}=E_{1},$ so that one can replace $\langle H_{0}\rangle$ by $\langle H_{0}\rangle-E_{1}$ in the denominator of Eq. (15).
One might wonder under which conditions the QSL bound is attained. To simplify the discussion, we consider the Hamiltonian $H_{0}^{\prime}=H_{0}-E_{1}$ with a ground state of zero energy, and the associated quantity $D(t)$. All the terms in the sum of Eq. (13) must saturate the inequality Eq. (14). The first term, by construction, corresponds to $x=0$, and there can be at most one additional term corresponding to $x=(E_{j}-E_{1})\tau/\hbar=\pi$ for some $j>1$. The only way to saturate the bound is therefore to choose the evolution time as $\tau=\pi/\hbar(E_{j}-E_{1}),$ and to have a Hamiltonian of the form:
$H_{0}=\sum_{n\in\mathcal{S}_{1}}E_{1}|n\rangle\langle n|+\sum_{n\in\mathcal{S}_{j}}E_{j}|n\rangle\langle n|$ (16)
@@
The coincidence of the MT and ML QSL bounds occurrs when $\Delta H=\langle H\rangle-E_{1}$. Since $\langle H\rangle-E_{1}=\frac{n}{d}(E_{j}-E_{1})$ and $\Delta H=\frac{1}{d}\sqrt{n(d-n)}(E_{j}-E_{1})$ (with $n={\rm card}\mathcal{S}_{j}$), this is simply equivalent to $n=d/2$. The two bounds coincide and saturate when the quantum system exhibits dynamics essentially identical to that of a two-level system.

### v2

@@ Abstract
We present a geometric optimization method for implementing quantum gates by optimally controlling the Hamiltonian parameters, with the goal of approaching the Mandelstam-Tamm Quantum Speed Limit (MT-QSL). Achieving this bound requires satisfying precise geometric conditions that govern the evolution of quantum states. We extend this geometric framework to quantum unitary operators in arbitrary dimensions and analyze the conditions necessary for saturation of the bound. Additionally, we show that the quantum brachistochrone, when generalized to operators, does not, in general, saturate the MT-QSL bound. Finally, we propose a systematic optimal control strategy based on geometric principles to approach the quantum speed limit for unitary driving. We illustrate this optimization method on a set of four well-known two-qubit quantum gates. Our procedure significantly reduces the deviation from the optimal quantum speed limit while preserving high quantum fidelity.
@@
Approaching optimality in quantum driving is both a challenging and crucial task for quantum information processing. Reducing the evolution time while keeping resources constant enhances efficiency and can increase the resilience of quantum protocols to external perturbations. The concept of Quantum Speed Limit (QSL) provides a quantitative bound for this requirement. The QSL literature commonly presents two main criteria, depending on the specific problem: a global criterion that defines the minimal time for an initial state to reach an orthogonal subspace, and a local criterion that bounds the instantaneous angular separation between the initial and evolving states. This concept has been a subject of growing interest, leading to numerous extensions [1, 2, 3, 4, 5, 6, 7, 8, 9], with the recent generalization to open  [10, 11, 12, 13], non-Markovian [14] and non-linear [15, 16] quantum systems. Among recent developments, we highlight efforts to establish unified geometric definitions of the QSL [17], the investigation of the relation of the QSL with the system quantumness [18, 19], with interference effects [20, 21], or with statistical properties [22, 23] of the system, and finally an effort to approach the QSL in concrete experimental setups [24, 25]. The QSL is fundamentally linked to the geometry of quantum state space [26] and involves identifying quantum trajectories that correspond to geodesics in the Fubini-Study metric. When formulated as a quantum brachistochrone problem, it allows for analytical [27, 28] or approximate [29] solutions for specific classes of Hamiltonians. However, optimal solutions to the quantum brachistochrone problem do not always saturate the QSL bounds associated with the evolution of a single quantum state [30]. While a significant body of literature has focused on QSLs for the evolution of individual states, the exploration of QSLs governing the evolution of sets of quantum states remains in its early stages [31, 32, 33, 34, 35, 36].
@@
The goal of this paper is to present an alternative derivation of such bounds, inspired by the approach of L. Vaidman [3], which also introduces a geometric optimization method for implementing time-optimal unitary quantum gates. More specifically, we identify two necessary criteria for optimal driving: one involves a phase condition, while the other has a geometric interpretation. Indeed, the deviation from the optimal Mandelstam–Tamm QSL bound grows as the evolution operator departs from a specific time-dependent plane in the Hilbert space of quantum operators.
@@
In Sec. I, we generalize the Mandelstam-Tamm bound on the quantum speed limit using an approach similar to that developed by L. Vaidman [3]. We then discuss the generalization of the Margolus-Levitin bound for time evolution operators. In Sec. II, we explore the conditions under which the bound is saturated, illustrating these with various examples, and analyze the quantum brachistochrone from the QSL perspective. In Sec. III, we detail the quantum optimal driving procedure designed to achieve the maximal Quantum Speed Limit, based on Pontryagin’s maximum principle and using geometrically derived cost functionals. Finally, in Sec. IV, we illustrate our protocol for various unitary gates in a four-dimensional space. We show the importance of the control over a large number of Hamiltonian parameters, compare the efficiency of the algorithm with respect to different cost functionals, and discuss the results of another optimization scheme.
@@ I Quantum Speed Limits for unitary operators
[…]
@@ I.1 Mandelstamm and Tamm bound for quantum operators
Using geometric arguments, L. Vaidman derived the minimal time required to evolve a given initial quantum state into an orthogonal one  [3]. Following a similar approach, we derive a QSL for the full quantum evolution operator and establish a lower bound on the evolution time required to reach a target quantum gate $U_{\rm target}$.
@@
where $U_{\perp}(t)$ is a linear operator such that $\langle U(t),U_{\perp}(t)\rangle=0$ and $\langle U_{\perp}(t),U_{\perp}(t)\rangle=1$ (note that $U_{\perp}(t)$ is in general non-unitary). Projecting both sides of the equation onto the operator $U(t)$, we obtain $\lambda_{/\!/}(t)=\frac{1}{d}{\rm Tr}[U^{\dagger}(t)H(t)U(t)]=\frac{1}{d}{\rm Tr}[H(t)]=\langle H(t)\rangle$ - where we define $\langle O\rangle=\frac{1}{d}{\rm Tr}[O]$ for any operator $O$. Next, we take the scalar product with the Hermitian conjugate $H(t)U(t)$ on both sides of Eq. (1), and readily obtain $\frac{1}{d}{\rm Tr}[H(t)^{2}]=|\lambda_{/\!/}(t)|^{2}+|\lambda_{\perp}(t)|^{2}$ where we have used $H(t)^{\dagger}=H(t)$, $\langle U(t),U(t)\rangle=\langle U_{\perp}(t),U_{\perp}(t)\rangle=1$ and $\langle U_{\perp}(t),U(t)\rangle=\langle U(t),U_{\perp}(t)\rangle=0.$ As a result, we find $|\lambda_{\perp}(t)|^{2}=\langle H(t)^{2}\rangle-\langle H(t)\rangle^{2}$, i.e. $|\lambda_{\perp}(t)|=\Delta H(t).$ Consequently, the operator $U_{\perp}$ can be expressed as
$U_{\perp}(t)=\frac{1}{\Delta H(t)}\left(H(t)U(t)-\langle H(t)\rangle U(t)\right)$ (2)
@@
$\frac{d|\langle U_{0},U(t)\rangle|^{2}}{dt}=\frac{-2\Delta H(t)}{\hbar}{\rm Re}\left[i\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle\right]$
We have omitted the term proportional to the average energy, which corresponds to a purely imaginary number. The initial evolution operator $U_{0}$ can be decomposed as
@@
[…]
@@
$\left|\frac{d|\langle U_{0},U(t)\rangle|^{2}}{dt}\right|\leq\frac{2\Delta H(t)}{\hbar}|\langle U_{0},U(t)\rangle|\sqrt{1-|\langle U_{0},U(t)\rangle|^{2}}.$ (8)
By setting $\cos\phi=|\langle U_{0},U(t)\rangle|$, the inequality (8) can be expressed as an upper bound on the time derivative of the angle $\phi$:
$|\dot{\phi}|\leq\frac{\Delta H}{\hbar}.$ (9)
By construction, the initial angle is $\phi(0)=0$. For a time-independent Hamiltonian $H$, the minimal time $\tau$ to reach a given target quantum gate $U_{\rm target}$ is given by:
$\tau=\frac{\hbar\phi_{\rm target}}{\Delta H}\quad\text{where}\quad\phi_{\rm target}={\rm Arccos}\left[\frac{|{\rm Tr}[U_{\rm target}]|}{d}\right].$ (10)
For a target gate with vanishing trace, $\phi_{\rm target}=\pi/2$. In this specific case, we retrieve a result analogous to the Mandelstamm-Tamm time bound for reaching an orthogonal quantum state.
Saturation of this QSL requires that two conditions be simultaneously fulfilled
$|\langle U_{0},U(t)\rangle|^{2}+|\langle U_{0},U_{\perp}(t)\rangle|^{2}$ $=$ $1$
${\rm Re}\left[\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle\right]$ $=$ $0$ (11)
The first condition states that $U_{\perp\perp}(t)=0,$ meaning that the time derivative $\dot{U}(t)$ must always lie in the plane spanned by $\{U_{0},U(t)\}$. Since the evolution operator $U(t)$ must remain unitary at all times, this imposes strong constraints on the possible quantum paths saturating the QSL bound. In practice, it is not guaranteed that such a path exists in Hilbert spaces with dimensions larger than two.
In such cases, one can quantify the deviation from the QSL when these conditions are not satisfied. Using $U_{\perp\perp}(t)||=\sqrt{\sin^{2}\phi(t)-|\langle U_{0},U_{\perp}(t)\rangle|^{2}}$ and $\beta(t)={\rm Arg}[\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle]$, one has
[…]
The ideal case of QSL saturation corresponds to $U_{\perp\perp}(t)||=0$ (plane condition) and to $\beta(t)=\pm\frac{\pi}{2}$ (phase condition) at all times. An optimization method based on these two relations will be presented in Sec. III.
@@ I.2 Margolus-Levitin quantum speed limit
the Margolus–Levitin bound  [37] can also be recovered within this operatorial formalism. To this end, we consider a time-independent Hamiltonian $H_{0}=\sum_{n=1}^{d}E_{n}|n\rangle\langle n|$ with $E_{1}\leq E_{2}\leq...\leq E_{d}$ possibly with degenerate eigenvalues. The associated evolution operator reads $U(t)=\sum_{n=1}^{d}e^{-\frac{i}{\hbar}E_{n}t}|n\rangle\langle n|,$ and one finds
$D(t)=\frac{1}{d}{\rm Tr}\left[U_{0}^{\dagger}U(t)\right]=\frac{1}{d}\sum_{n=1}^{d}e^{-\frac{i}{\hbar}E_{n}t},$ (13)
where we take the initial evolution operator as $U_{0}=\mathbb{I}$ the same reasoning applies to any other gate $U_{f}$). One can proceed in the same way as in the proof for single states (with here $|c_{n}|^{2}\equiv 1/d$). Thus, we have that
${\rm Re}\left[D(t)\right]$ $=$ $\frac{1}{d}\sum_{n=1}^{d}\cos\left(\frac{E_{n}}{\hbar}t\right)$
$\geq$ $\frac{1}{d}\sum_{n}\left[1-\frac{2}{\pi}\left(\frac{E_{n}t}{\hbar}+\sin\left(\frac{E_{n}t}{\hbar}\right)\right)\right],$
@@
valid for $x\geq 0$, which becomes an equality if and only if $x=0,\pi$. Finally, the right-hand side of the inequality (I.2) can be expressed in terms of the average energy and imaginary part of $D(t)$, so that:
${\rm Re}\left[D(t)\right]\geq 1-\frac{2}{\pi}\langle H_{0}\rangle\frac{t}{\hbar}+\frac{2}{\pi}{\rm Im}[D(t)].$ (16)
This can be interpreted in terms of minimum time $\tau$ to reach a given target quantum gate $U_{\rm target}$:
$\tau\geq\frac{\hbar\pi}{2\langle H_{0}\rangle}\left(1-D^{R}_{\rm target}+\frac{2}{\pi}D^{I}_{\rm target}\right),$ (17)
where we have noted $D^{R}_{\rm target}=\frac{1}{d}{\rm Re}\left[{\rm Tr}[U_{\rm target}]\right]$ and $D^{I}_{\rm target}=\frac{1}{d}{\rm Im}\left[{\rm Tr}[U_{\rm target}]\right]$. For the particular case where the target quantum gate has a null trace, one obtains
$\tau\geq\frac{\pi\hbar}{2\langle H_{0}\rangle},$ (18)
which is the quantum-operator equivalent of the Margolus-Levitin bound for individual quantum states.
When ${\rm Tr}[U_{\rm target}]=0$, a tighter bound can be derived. The argument is valid for any constant and positive Hamiltonian $H_{0}$. The replacement $H_{0}^{\prime}=H_{0}-E^{\prime}$ merely adds a global phase factor and thus leaves invariant the time $\tau$ such that $D(\tau)=0$. As long as $E^{\prime}\leq E_{1}$ to guarantee the positivity of the Hamiltonian, the argument above yields a similar bound with $\langle H_{0}^{\prime}\rangle=\langle H_{0}\rangle-E^{\prime}$ instead of $\langle H_{0}\rangle.$ The tighter bound is obtained with $E^{\prime}=E_{1},$ so that one can replace $\langle H_{0}\rangle$ by $\langle H_{0}\rangle-E_{1}$ in the denominator of Eq. (18). When ${\rm Tr}[U_{\rm target}]\neq 0$, adding a constant energy term can alter the time $\tau$ required to achieve the quantum gate. Therefore, the denominator in Eq. (17) must remain $\langle H_{0}\rangle$.
To saturate the Margolus-Levitin bound, the only admissible values of $x=E_{n}\tau$ in inequality (I.2) are $\{0,\pi\}$. Consequently, only quantum gates with a real-valued trace can be attained through a path that saturates the Margolus–Levitin bound.
Let’s reconsider the case of a target gate where $D(\tau)=\frac{1}{d}{\rm Tr}[U_{\rm target}]=0$. One might wonder under which conditions the QSL bound is attained. To simplify the discussion, we consider a Hamiltonian $H_{0}^{\prime}=H_{0}-E_{1}$ with a ground state of zero energy, and the associated quantity $D(t)$. All the terms in the sum of Eq. (I.2) must saturate the inequality Eq. (22). The first term, by construction, corresponds to $x=0$, and there can be at most one additional term corresponding to $x=(E_{j}-E_{1})\tau/\hbar=\pi$ for some $j>1$. The only way to saturate the bound is therefore to choose the evolution time as $\tau=\pi/\hbar(E_{j}-E_{1}),$ and to have a Hamiltonian of the form:
$H_{0}=\sum_{n\in\mathcal{S}_{1}}E_{1}|n\rangle\langle n|+\sum_{n\in\mathcal{S}_{j}}E_{j}|n\rangle\langle n|,$ (19)
@@
The coincidence of the MT and ML QSL bounds occurs when $\Delta H=\langle H\rangle-E_{1}$. Since $\langle H\rangle-E_{1}=\frac{n}{d}(E_{j}-E_{1})$ and $\Delta H=\frac{1}{d}\sqrt{n(d-n)}(E_{j}-E_{1})$ (with $n={\rm card}\mathcal{S}_{j}$), this reduces to the simple condition $n=d/2$. In this case, the two bounds coincide and are saturated when the quantum system exhibits dynamics essentially identical to those of a two-level system.